### v2

@@ 3.1 Mitigating Heterogeneity through Relevancy Scoring
As shown in Table 3.2, in both case studies, the relevancy scores from LLMs indicated higher agreement, as measured by Gwet’s AC1 and Fleiss’ kappa, than the dichotomous labels returned by the LLMs. This result suggests that (1) the scoring method can effectively mitigate the heterogeneity in the labeling decisions made by multiple LLMs, and (2) the reasoning of LLMs exhibited higher similarity than their binary classifications.
@@ 3.2 Enhancing Accuracy through Ensembling
In case study 2, as shown in Figure 4, LLMs exhibited varying capabilities in determining the presence of weight stigma in ED patients’ narratives. Qwen2.5-7B-Instruct was too conservative in labeling positive cases, resulting in low sensitivity and high precision or specificity, whereas Vicuna-7b-v1.5 was too generous, resulting in high sensitivity but low precision or specificity. Nevertheless, the ensemble of three or more LLMs achieved higher F1-score, precision, specificity and sensitivity than individual LLMs when classifying human annotations. Similarly, the ensembles of scores showed the highest AUPRC compared to individual LLMs. The removal of the detected outlier LLM, Qwen2.5-7B-Instruct, led to increased sensitivity but decreased precision and specificity. The overall F1-score remained the same as that of the ensemble of four LLMs. Note that specificity was included in the evaluation for this case study, due to the relatively balanced prevalence of human-annotated weight stigma–70% among 1080 posts.
{tableorg}
Evaluation of agreement on labeling and scoring1 conducted with four open-source, small-sized language models2. 0pt Gwet’s AC1 [95% CI] Fleiss’ kappa [95% CI] Topic Labels Scores Labels Scores binge 0.939 [0.937, 0.940] 0.924 [0.922, 0.925] 0.651 [0.644, 0.659] 0.702 [0.695, 0.709] bodyhate 0.827 [0.825, 0.830] 0.852 [0.850, 0.854] 0.451 [0.445, 0.457] 0.484 [0.477, 0.492] calorie 0.900 [0.898, 0.902] 0.881 [0.879, 0.883] 0.671 [0.665, 0.676] 0.722 [0.718, 0.727] crave 0.858 [0.856, 0.860] 0.884 [0.882, 0.886] 0.354 [0.347, 0.361] 0.408 [0.398, 0.418] depressedmood 0.774 [0.771, 0.777] 0.854 [0.852, 0.856] 0.321 [0.316, 0.325] 0.442 [0.435, 0.450] ed 0.902 [0.900, 0.904] 0.879 [0.877, 0.881] 0.689 [0.683, 0.694] 0.758 [0.753, 0.763] exercise 0.832 [0.829, 0.835] 0.781 [0.779, 0.784] 0.730 [0.726, 0.734] 0.647 [0.643, 0.651] fearfood 0.946 [0.944, 0.947] 0.946 [0.944, 0.947] 0.300 [0.291, 0.309] 0.354 [0.340, 0.369] feargain 0.904 [0.902, 0.905] 0.903 [0.901, 0.905] 0.377 [0.369, 0.385] 0.360 [0.351, 0.369] gain 0.880 [0.878, 0.882] 0.864 [0.862, 0.866] 0.563 [0.557, 0.568] 0.573 [0.567, 0.579] idealbody 0.790 [0.787, 0.792] 0.828 [0.825, 0.830] 0.246 [0.240, 0.251] 0.232 [0.225, 0.238] loss 0.751 [0.748, 0.754] 0.772 [0.769, 0.774] 0.546 [0.542, 0.551] 0.596 [0.591, 0.601] protein 0.796 [0.793, 0.799] 0.820 [0.818, 0.823] 0.452 [0.447, 0.458] 0.596 [0.589, 0.603] relation 0.338 [0.333, 0.342] 0.825 [0.823, 0.827] 0.057 [0.053, 0.061] 0.350 [0.343, 0.357] restrict 0.761 [0.757, 0.764] 0.793 [0.791, 0.796] 0.495 [0.490, 0.500] 0.526 [0.521, 0.532] Data 1: ED and Dieting Reddit Posts 0pt Gwet’s AC1 [95% CI] Fleiss’ kappa [95% CI] Topic Labels Scores Labels Scores weightstigma 0.601 [0.572, 0.629] 0.597 [0.575, 0.618] 0.187 [0.152, 0.222] 0.300 [0.269, 0.330] Data 2 (Protected): ED Patients Experiences
1 Embedding scores were calculated using all-mpnet-base-v2 and binned into 10 ordinal levels for agreement evaluation. 2 The four open- source, small-sized language models are Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct, Mistral-7B-Instruct-v0.3, and Vicuna-7b-v1.5.
Figure 2: Performance in predicting human labels of 15 topics related to eating disorders and dieting in Reddit posts, using four open-source, locally-deployable LLMs and all combinations of their ensembles. Panel A illustrates the performance of using relevancy scores or labels returned by LLMs and their ensembles to classify human annotations. The AUPRC metric evaluates relevancy scores, while F1-score, precision, and sensitivity evaluate performance based on labels. Different colors represent various LLMs and their ensembles. Performance metrics for GPT-4o are indicated by horizontal dashed lines. For ensembles of two and three LLMs, error bars indicate the minimum, maximum, and the 25th and 75th percentiles across all combinations of the respective ensemble size. Panel B illustrates the performance increases achieved by using ensembles of LLMs to classify human annotations, compared to the median performance of the individual LLMs included in each ensemble.
@@
Figure 4: Performance in predicting human annotations of weight stigma experienced by ED patients during treatment, using four open-source, locally-deployable LLMs and all combinations of their ensembles. Panel A illustrates the performance of using relevancy scores or labels returned by LLMs and their ensembles to classify human annotations. The AUPRC metric evaluates relevancy scores, while F1-score, precision, and sensitivity evaluate performance based on labels. Different colors represent various LLMs and their ensembles. For ensembles of two and three LLMs, error bars indicate the minimum, maximum, and the 25th and 75th percentiles across all combinations of the respective ensemble size. Panel B illustrates the performance increases achieved by using ensembles of LLMs to classify human annotations, compared to the median performance of the individual LLMs included in each ensemble.
@@ 4 Discussion
In psychology research, annotating granular topics, such as pathological behaviors and mentalities in free-form texts like patient narratives, is typically labor-intensive and time-consuming. Without user consent for external use, LLM-assisted labeling techniques that rely on closed-source models through APIs can hardly be applied to such free-text data. These constraints also limit the use of large-sized LLMs ($>$100 billion parameters) because they cannot be deployed locally in most research settings.
In this study, we proposed an ensembling framework that leverages the heterogeneity of diverse, open-source, locally-deployable LLMs for topic labeling and relevancy scoring on free-text data. This ensembling is guided by a novel relevancy scoring methodology that utilizes embedding distances between the topic description and LLMs’ reasoning. The topic labeling and scoring by multiple LLMs mirror the human labeling process performed by several human coders. The main idea of ensembling labels from multiple LLMs is to eliminate potential false positives from the union of positive labels returned by multiple LLMs, which inherits all false positives from each LLM. Conversely, the intersection of positive labels, based on over-half agreement among LLMs, is likely to exclude more true positives. Thus, label ensembling seeks a balance between the union and the intersection, guided by the ensemble of relevancy scores.
The framework was not designed to replace human evaluation, but rather to alleviate the burden of the human labeling process by providing a preliminary label and a relevancy score for each text. The final labeling decision, typically made by domain experts, can reference these preliminary labels, focusing particularly on positive labels with low relevancy scores (potential false positives) and negative labels with positive scores (potential false negatives).
@@
    The ensemble of LLMs effectively leveraged the heterogeneity of the LLMs, achieving the highest accuracy in predicting human annotations. In the first case study, the ensemble of LLMs also demonstrated the highest accuracy in predicting closed-source large-sized GPT-4o’s labels. The improvement in accuracy was evidenced by superior F1-score, superior precision, and sensitivity compared to individual LLMs.
@@
    The ensemble of a larger number of LLMs demonstrated more reliable performance improvements and a more optimal precision-recall trade-off. In contrast, ensembles of only two LLMs appeared more vulnerable to the variability in individual LLM performance.
  4. 4.

### v4

@@ 3 Results
Table 3: Evaluation of agreement on labeling and scoring1 conducted with four open-source, small-sized language models2. Bold values indicate higher agreement in scores than in labels returned by language models.
Gwet’s AC1 [95% CI] Fleiss’ kappa [95% CI]
Topic Labels Scores Labels Scores
binge 0.939 [0.937, 0.940] 0.924 [0.922, 0.925] 0.651 [0.644, 0.659] 0.702 [0.695, 0.709]
bodyhate 0.827 [0.825, 0.830] 0.852 [0.850, 0.854] 0.451 [0.445, 0.457] 0.484 [0.477, 0.492]
calorie 0.900 [0.898, 0.902] 0.881 [0.879, 0.883] 0.671 [0.665, 0.676] 0.722 [0.718, 0.727]
crave 0.858 [0.856, 0.860] 0.884 [0.882, 0.886] 0.354 [0.347, 0.361] 0.408 [0.398, 0.418]
depressedmood 0.774 [0.771, 0.777] 0.854 [0.852, 0.856] 0.321 [0.316, 0.325] 0.442 [0.435, 0.450]
ed 0.902 [0.900, 0.904] 0.879 [0.877, 0.881] 0.689 [0.683, 0.694] 0.758 [0.753, 0.763]
exercise 0.832 [0.829, 0.835] 0.781 [0.779, 0.784] 0.730 [0.726, 0.734] 0.647 [0.643, 0.651]
fearfood 0.946 [0.944, 0.947] 0.946 [0.944, 0.947] 0.300 [0.291, 0.309] 0.354 [0.340, 0.369]
feargain 0.904 [0.902, 0.905] 0.903 [0.901, 0.905] 0.377 [0.369, 0.385] 0.360 [0.351, 0.369]
gain 0.880 [0.878, 0.882] 0.864 [0.862, 0.866] 0.563 [0.557, 0.568] 0.573 [0.567, 0.579]
idealbody 0.790 [0.787, 0.792] 0.828 [0.825, 0.830] 0.246 [0.240, 0.251] 0.232 [0.225, 0.238]
loss 0.751 [0.748, 0.754] 0.772 [0.769, 0.774] 0.546 [0.542, 0.551] 0.596 [0.591, 0.601]
protein 0.796 [0.793, 0.799] 0.820 [0.818, 0.823] 0.452 [0.447, 0.458] 0.596 [0.589, 0.603]
relation 0.338 [0.333, 0.342] 0.825 [0.823, 0.827] 0.057 [0.053, 0.061] 0.350 [0.343, 0.357]
restrict 0.761 [0.757, 0.764] 0.793 [0.791, 0.796] 0.495 [0.490, 0.500] 0.526 [0.521, 0.532]
(a) Data 1: ED and Dieting Reddit Posts
Gwet’s AC1 [95% CI] Fleiss’ kappa [95% CI]
Topic Labels Scores Labels Scores
weightstigma 0.601 [0.572, 0.629] 0.597 [0.575, 0.618] 0.187 [0.152, 0.222] 0.300 [0.269, 0.330]
(b) Data 2 (Protected): ED Patients Experiences
1 Embedding scores were calculated using all-mpnet-base-v2 and binned into 10 ordinal levels for agreement evaluation.
2 The four open- source, small-sized language models are Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct, Mistral-7B-Instruct-v0.3, and Vicuna-7b-v1.5.
@@ 3.1 Mitigating Heterogeneity through Relevancy Scoring
As shown in Table 3, in both case studies, the relevancy scores from LLMs demonstrated higher agreement for most topics, as measured by Gwet’s AC1 and Fleiss’ kappa, than the dichotomous labels returned by the LLMs. This result suggests that (1) the scoring method can effectively mitigate the heterogeneity in the labeling decisions made by multiple LLMs, and (2) the reasoning of LLMs exhibited higher similarity than their binary classifications. Gwet’s AC1 and Fleiss’s Kappa differ because they calculate chance agreement differently: Gwet’s AC1 uses an adjusted chance agreement to account for imbalances in category prevalence, whereas Fleiss’s Kappa relies directly on marginal distributions.
@@ 3.2 Enhancing Accuracy through Ensembling
In case study 2, as shown in Figure 6, LLMs exhibited varying capabilities in determining the presence of weight stigma in ED patients’ narratives. Qwen2.5-7B-Instruct was too conservative in labeling positive cases, resulting in low sensitivity and high precision, whereas Vicuna-7b-v1.5 was too generous, resulting in high sensitivity but low precision. Nevertheless, the ensemble of three or more LLMs achieved higher F1-score, precision, and sensitivity than individual LLMs when classifying human annotations. Similarly, the ensembles of scores showed the highest AUPRC compared to individual LLMs. The removal of the detected outlier LLM, Qwen2.5-7B-Instruct, led to increased sensitivity but slightly decreased precision. The overall F1-score remained the same as that of the ensemble of four LLMs.
Figure 4: Performance in predicting human labels of 15 topics related to eating disorders and dieting in Reddit posts, using four open-source, locally-deployable LLMs and all combinations of their ensembles. Panel A. illustrates the performance of using relevancy scores or labels returned by LLMs and their ensembles to classify human annotations. The AUPRC metric evaluates relevancy scores, while F1-score, precision, and sensitivity evaluate performance based on labels. Different colors represent various LLMs and their ensembles. Performance metrics for GPT-4o are indicated by horizontal dashed lines. For ensembles of two and three LLMs, error bars indicate the minimum, maximum, and the 25th and 75th percentiles across all combinations of the respective ensemble size. Panel B. illustrates the performance increases achieved by using ensembles of LLMs to classify human annotations, compared to the median performance of the individual LLMs included in each ensemble.
@@
Figure 6: Performance in predicting human annotations of weight stigma experienced by ED patients during treatment, using four open-source, locally-deployable LLMs and all combinations of their ensembles. Panel A. illustrates the performance of using relevancy scores or labels returned by LLMs and their ensembles to classify human annotations. The AUPRC metric evaluates relevancy scores, while F1-score, precision, and sensitivity evaluate performance based on labels. Different colors represent various LLMs and their ensembles. For ensembles of two and three LLMs, error bars indicate the minimum, maximum, and the 25th and 75th percentiles across all combinations of the respective ensemble size. Panel B. illustrates the performance increases achieved by using ensembles of LLMs to classify human annotations, compared to the median performance of the individual LLMs included in each ensemble.
@@ 4 Discussion
In psychology research, annotating granular topics, such as pathological behaviors and mentalities in free-form texts like patient narratives, is typically labor-intensive and time-consuming. Without user consent for external use, LLM-assisted labeling techniques that rely on closed-source models through APIs can hardly be applied to such free-text data. These constraints also limit the use of large-sized LLMs (>>100 billion parameters) because they cannot be deployed locally in most research settings.
In this study, we proposed an ensembling framework that leverages the heterogeneity of diverse, open-source, locally-deployable LLMs for topic labeling and relevancy scoring on free-text data. This ensembling is guided by a relevancy scoring methodology that utilizes embedding distances between the topic description and LLMs’ reasoning. The topic labeling and scoring by multiple LLMs mirror the human labeling process performed by several human coders. The main idea of ensembling labels from multiple LLMs is to eliminate potential false positives from the union of positive labels returned by multiple LLMs, which inherits all false positives from each LLM. Conversely, the intersection of positive labels, based on majority voting among LLMs, may be closer to the ground truth but is likely to exclude true positives. Thus, our label ensembling seeks a balance between the union and the intersection, guided by the ensemble of relevancy scores. The framework was not designed to replace human evaluation, but rather to alleviate the burden of the human labeling process by providing a preliminary label and a relevancy score for each text. The final labeling decision, typically made by domain experts, can reference these preliminary labels, focusing particularly on positive labels with low relevancy scores (potential false positives) and negative labels with positive scores (potential false negatives).
@@
    The ensemble of LLMs effectively leveraged the heterogeneity of the LLMs, achieving the highest accuracy in predicting human annotations. In the first case study, the ensemble of LLMs also demonstrated the highest accuracy in predicting closed-source large-sized GPT-4o’s labels. The improvement in accuracy was evidenced by superior F1-score, superior precision, and sensitivity compared to individual LLMs. Moreover, the ensemble of a larger number of LLMs demonstrated more reliable performance improvements and a more optimal precision-recall trade-off. In contrast, ensembles of only two LLMs appeared more vulnerable to the variability in individual LLM performance.
@@
We also observed heterogeneity in the performance of LLMs when annotating and rating different topics. Although the ensembles generally outperform individual LLMs, for some topics, all LLMs exhibited worse performance, as measured by F1, precision, and recall. This might be due to the complexity of the topic definitions. For example, topics such as ’fear of certain foods’ and ’fear of weight gain’ posed challenges not only for open-source small-sized LLMs but also for GPT-4o to detect. These topics required understanding both the emotional content, ’fear,’ and the context of food or weight gain from the text. Similarly, the definition of ’ideal body image’ is complex; it may involve being muscular and having lean body mass for males, while for females, it often pertains to being thin and skinny. Although the ensembles maintained higher performances than individual LLMs, enhancing LLMs’ understanding of topic definitions through more complex prompting techniques, such as chain of thought prompting [72] or soft prompting [98], or fine-tuning on datasets annotated with these topics, are important next steps.
@@
Appendix A Inter-rater agreement coefficients
A.1 Gwet’s AC1
Gwet’s AC1 [81] is an inter-rater agreement coefficient designed to overcome the limitations of Cohen’s or Fleiss’ Kappa, particularly their sensitivity to prevalence and marginal distribution. Like Kappa, AC1 is defined as the ratio of observed agreement beyond chance over the maximum possible agreement beyond chance:
$AC1=\frac{P_{o}-P_{e}^{*}}{1-P_{e}^{*}},$
where $P_{o}$ is the observed proportion of agreement among raters, and $P_{e}^{*}$ is the expected agreement by chance.
Let $n_{ij}$ denote the number of raters who assigned item $i$ to category $j$, with $n$ raters, $N$ items and $k$ categories in total. The observed agreement $P_{o}$ is given by
$P_{o}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{k}\frac{n_{ij}(n_{ij}-1)}{n(n-1)},$
Let $p_{j}=\frac{1}{N}\sum_{i=1}^{N}\frac{n_{ij}}{n}$ be the overall proportion of ratings assigned to category $j$. The adjusted chance agreement $P_{e}^{*}$ is given by
$P_{e}^{*}=\frac{1}{k-1}\sum_{j=1}^{k}p_{j}(1-p_{j}).$
This adjustment prevents artificially low agreement coefficients when one category is dominant, making AC1 more robust in cases of imbalanced class distributions. For comparison, Fleiss’s Kappa is defined in the next section.
A.2 Fleiss’s Kappa
Fleiss’s Kappa [83, 99] is a generalization of Cohen’s Kappa for assessing inter-rater agreement among $n\geq 2$ raters when assigning categorical ratings to $N$ items. Like other chance-corrected coefficients, it is defined as
$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}},$
where $P_{o}$ is the observed agreement and $P_{e}$ is the expected agreement by chance.
Let $n_{ij}$ denote the number of raters who assigned item $i$ to category $j$, with $n$ raters, $N$ items and $k$ categories in total. The observed agreement $P_{o}$ across all items is given by
$P_{o}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{k}\frac{n_{ij}(n_{ij}-1)}{n(n-1)}.$
The chance agreement $P_{e}$ is given by
$P_{e}=\sum_{j=1}^{k}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{n_{ij}}{n}\right)^{2}.$
Although widely used, the chance agreement $P_{e}$ is sensitive to prevalence and marginal distributions, which can result in the “high agreement, low reliability” paradox when the target category has low prevalence.